OpenStudy (anonymous):

Carl conducted an experiment to determine if there is a difference in the mean body temperature between men and women. He found that the mean body temperature for a sample of 100 men was 91.1 with a population standard deviation of 0.52 and the mean body temperature for a sample of 100 women was 97.6 with a population standard deviation of 0.45. Assuming the population of body temperatures for men and women is normally distributed, calculate the 98% confidence interval and the margin of error for the mean body temperature for both men and women.

OpenStudy (anonymous):

Using complete sentences, explain what these confidence intervals mean in the context of the problem.

OpenStudy (misssunshinexxoxo):

Men's CI: x-bar = 91.1 ME = 2.3263*0.52 = 1.210 98% CI:: 91.1-1.21 < u < 91.1+1.21 98% CI: 89.89 < u < 92.31

OpenStudy (misssunshinexxoxo):

Women's CI: x-bar = 97.6 ME = 2.3263*0.45 = 1.05 98% CI: 97.6-1.05 < u < 97.6+1.05 98% CI: 96.55 < u < 98.65

OpenStudy (misssunshinexxoxo):

"He found that the mean body temperature for a sample of 100 men was 91.1 with a population standard deviation of 0.52 and the mean body temperature for a sample of 100 women was 97.6 with a population standard deviation of 0.45."

OpenStudy (anonymous):

and could u explain the variables again plz

OpenStudy (misssunshinexxoxo):

100 men was 91.1, deviation was 0.52, and mean sample was 100 women

OpenStudy (misssunshinexxoxo):

those are your variables

OpenStudy (anonymous):

ok thanks! what a lifesaver